OpenStudy (anonymous):

1) How would the graph of the function y = x2 + 4 be affected if the function were changed to y = x2 + 1? A The graph would shift 3 units up. B The graph would shift 3 units down. C The graph would shift 3 units to the right. D The graph would shift 3 units to the left. 2) What is the effect on the graph of the equation y = x2 + 1 when it is changed to y = x2 + 5? A The slope of the graph changes. B The curve translates in the positive x direction. C The graph is congruent, and the vertex of the graph moves up the y-axis. D The graph narrows.

OpenStudy (callisto):

Hints: For Q1: 1. Find the y-intercept of y=x^2+4 2. Find the y-intercept of y=x^2+1 3. Plot the two points (found in (1) and (2)) 4. See the difference.